OpenStudy (anonymous):

3/x^2-3x+4/x=1/x-3, please HELP

OpenStudy (lgbasallote):

multiply both sides by the LCM which is x^2(x-3) 3(x-3) - (3x+4)(x(x-3)) = x^2 3x - 9 - (3x+4)(x^2-3x) = x^2 3x - 9 - 3x^3 -9x^2 + 4x^2 -12x = x^2 -3x^3 -x^2 -5x^2 + 3x - 12x = 9 -3x^3 - 6x^2 - 9x = 9 -3x(x^2 - 2x - 3) = 9 x(x^2 - 2x -3) = -3 x = -3; (x-3)(x+1)=-3 so x = -3; x = 0; x= -4